OpenStudy (anonymous):

If a cylinder has a height of 12 feet and a radius of 3 feet, what is its surface area? volume? (Round your answers to two decimal places

OpenStudy (ranga):

Volume = (pi)r^2h Lateral Surface Area = 2(pi)rh Top and bottom surface area = (pi)r^2 (each) Add all the surface areas to get the total surface area of the cylinder.

OpenStudy (ranga):

Total surface area = 2(pi)rh + 2(pi)r^2